OpenStudy (anonymous):

Make sure you read the directions for each part carefully. This system of equations will be used in all three parts of this question. x-2y=0 5x+2y=24 Part 1: Solve the system by graphing. Draw the graphs on paper and indicate their intersection point. Type the ordered pair solution here as your answer to Part 1. Part 2: Solve the system using the substitution method. Show all work here and indicate the solution for the system as an ordered pair. Part 3: Solve the system using the addition method. Show all work here and indicate the solution for the system as an ordered pair.

OpenStudy (anonymous):

x=2y 10y+2y=24 y=24/12=2 x=4 (4,2)

OpenStudy (anonymous):

adding both, y is cancelled out 6x=24 x=4 y=2 (4,2)

OpenStudy (anonymous):

that website doest help really

OpenStudy (anonymous):

all it does is give u da answer but i need the explation to soooo

OpenStudy (anonymous):

*explantion

OpenStudy (anonymous):

graph cant be drawn over here. no whiteboard

OpenStudy (anonymous):

u can draw both the lines by taking their intercepts

OpenStudy (anonymous):

yea but the graph on that site is weird i dont get it

OpenStudy (anonymous):

x-2y=0 is the first line, x=2y, passes through origin n (!,2), got it?

OpenStudy (anonymous):

5x+2y=24 passes through (24/5,0) and(0,12)

OpenStudy (anonymous):

join the points to get thelines, the point of intersection gives the solution